A pound of tin is worth $2 more than a pound of copper. Eight pounds of tin are mixed with 2 pounds of copper to make bronze that sells for $4.50 per pound. How much is a pound of tin worth?
Answer by ankor@dixie-net.com(22740)

A pound of tin is worth $2 more than a pound of copper.
Eight pounds of tin are mixed with 2 pounds of copper to make bronze that sells for $4.50 per pound.
How much is a pound of tin worth?
:
let t = price of a l lb of tin
then
(t-2) = price of 1 lb of copper
:
Resulting amt = 10 lb
:
8t + 2(t-2) = 4.5(10)
8t + 2t - 4 = 45
10t = 45 + 4
10t = 49
t = 49/10
t = $4.90 price of a lb of tin
;
:
Check this )copper is worth 2.90 a lb
4.90(8) + 2.90(2) = 4.5(10)
39.20 + 5.80 = 45; confirms our solution of t = 4.90